# Improving $R(3,k)$ in just two bites

### v2

@@ Abstract
We present a random construction proving that the extreme off-diagonal Ramsey numbers satisfy $R(3,k)\geq\left(\frac{1}{2}+o(1)\right)\frac{k^{2}}{\log{k}}$ (conjectured to be asymptotically tight), improving the previously best bound $R(3,k)\geq\left(\frac{1}{3}+o(1)\right)\frac{k^{2}}{\log{k}}$. In contrast to all previous constructions achieving the correct order of magnitude, we do not use a nibble argument.
@@ 1 Introduction
The Ramsey number $R(\ell,k)$ is the least $N$ such that every $N$-vertex graph contains either a clique of size $\ell$ or an independent set of size $k$. In 1930 Ramsey [Ramsey1930] showed the existence of $R(\ell,k)$, and in the intervening century much effort has been dedicated to understanding their asymptotic behavior. The best current bounds on the diagonal Ramsey numbers $R(k,k)$ (which are known exactly only up to $k=4$, with $43\leq R(5,5)\leq 46$ and the upper bound only very recently announced by Angeltveit and McKay [angeltveit2025]) are
@@
for some $c>0$. The lower bound is one of the earliest instances of Erdős’ probabilistic method [Erdos47], and the upper bound was the first exponential improvement over a result of Erdős and Szekeres [ES35], only recently announced in 2023 by Campos, Griffiths, Morris and Sahasrabudhe [campos2025exponentialimprovementdiagonalramsey], subsequently optimized to a value of $c\approx 0.2$ (for sufficiently large $k$) by Gupta, Ndiaye, Norin and Wei [gupta2024optimizingcgmsupperbound].
The second most studied family of graph Ramsey numbers are the extreme off-diagonal numbers $R(\ell,k)$ for $\ell$ fixed and $k$ tending to infinity. The case $\ell=3$ yields the first non-trivial asymptotic question. The exciting history of the problem until the early 2000s is described by Spencer [Spencer2011] in a chapter of the book “Ramsey Theory” [SoiferRamsey], while the more recent history is very well captured by Campos, Jenssen, Michelen and Sahasrabudhe [campos2025]. We briefly summarize the results determining the order of magnitude. The upper bound of $R(3,k)=O(\frac{k^{2}}{\log{k}})$ goes back to work of Ajtai, Komlós and Szemeŕedi  [AjtaiKSz80, AjtaiKSz81] in the early 1980s, and in 1983 Shearer [Shearer83] reduced the implied constant to $R(3,k)\leq(1+o(1))\frac{k^{2}}{\log{k}}$. The lower bound was improved in several steps until Kim [Kim95] proved in 1995 that $R(3,k)\geq c\frac{k^{2}}{\log{k}}$, with a constant $c\approx\frac{1}{160}$. In his paper, which led to him winning the Fulkerson Prize, Kim uses a probabilistic method commonly known as a “nibble”, where small batches of edges are added randomly for many steps. This carefully avoids the appearance of triangles while keeping the edge distribution sufficiently random to destroy large independent sets. Spencer [Spencer2011] writes in 2011:
@@
Nevertheless, only a short time later in 2013, two groups announced that they had greatly improved on the constant in the lower bound, showing that $R(3,k)\geq\left(\frac{1}{4}+o(1)\right)\frac{k^{2}}{\log{k}}$. Bohman and Keevash [BohmanKeevash21] and Fiz Pontiveros, Griffiths and Morris [Fiz20] very carefully analyzed the triangle-free process, in which an edge, chosen uniformly at random from those which would not create a triangle, is added to the graph (and this operation is repeated until no such edges exist). Both of these papers go through serious technical efforts to control the randomness in the process, and in particular show that with high probability the triangle-free process terminates with independence number yielding this bound. Fiz Pontiveros, Griffiths and Morris went so far as to conjecture that $c=\frac{1}{4}$ is the correct constant for $R(3,k)$.
In another breakthrough 12 years later, Campos, Jenssen, Michelen and Sahasrabudhe [campos2025] showed that the story of the lower bound was not finished, and in fact $R(3,k)\geq\left(\frac{1}{3}+o(1)\right)\frac{k^{2}}{\log{k}}$, now conjecturing that $c=\frac{1}{2}$ should be the correct constant.
Conjecture 1.1 (Campos, Jenssen, Michelen and Sahasrabudhe [campos2025]).
@@
In this paper, we prove the lower bound in Conjecture 1.1.
@@
As all other previous results, we prove the following theorem concerning independence numbers, which easily implies Theorem 1.2.
@@
While inspired by ideas in [campos2025], the construction in Theorem 1.3 does not require a nibble, and thus circumvents many of the technical difficulties faced by the previous groups. Spencer states in the epilogue of his book chapter:
@@
If one were interested only in the correct order of magnitude, our construction could achieve Spencer’s dream, by combining slightly weaker proofs of the ingredients of Theorem 1.3 with Shearer’s one page proof of the upper bound.
Our method seems quite versatile and has already seen some applications to related problems. In Section 5 we give a direct application to hypergraph Ramsey numbers of stars, suggested to us by Mubayi. Most recently, Campos, Jenssen, Michelen and Sahasrabudhe together with the fourth author [campos2025polynomialimprovementoddcyclecomplete] have built on our method to find the first polynomial improvements over the edge deletion threshold for the cycle-complete Ramsey numbers $r(C_{\ell},K_{k})$, for all odd $\ell\geq 9$ as $k$ tends to infinity.
@@ 2 The Construction of $G$
Before formally presenting our construction witnessing Theorem 1.3, we provide some heuristic justification. The edge deletion method is a typical starting point in modern combinatorics when one wishes to construct a large graph $G$ which avoids a forbidden subgraph $H$, but carries certain other properties held by the random graph (in our case, $G$ should have no large independent sets). Implementing the method first requires one to calculate the edge deletion threshold $p_{H}$, so that (in expectation) the number of copies of $H$ in the random graph $G(n,p_{H})$ is of the same order as $p_{H}n^{2}$, the expected number of edges. If $p\ll p_{H}$ and one samples $G$ from $G(n,p)$ before deleting an edge from every copy of $H$, only $o(p_{H}n^{2})$ edges are lost. Some further effort can show that the resulting graph will have small independence number, because $G(n,p)$ had small independence number and only a small number of edges were removed. For many problems, this approach gives the best known constructions. To outperform such examples, one may try to construct an $H$-free graph with edge density larger than $p_{H}$, while still being careful to obtain some of the pseudorandomness that the random graph enjoys.
@@
Most recently, Campos, Jenssen, Michelen and Sahasrabudhe [campos2025] constructed a triangle-free graph with edge density $(1+o(1))\sqrt{\frac{2\log n}{3n}}$. To do so, they let $N=n/\log^{2}(n)$ and $p=\sqrt{\frac{\log(n)}{6n}}$ before sampling from $G_{1}$ from $G(N,p)$. Noting that $p=o(\sqrt{\frac{1}{N}})$ they can delete few edges to make the resulting graph (still on $N$ vertices) triangle free, before blowing up each vertex by a factor $\log^{2}(n)$ to obtain a new graph on $n$ vertices, still triangle-free and with edge density $\sqrt{\frac{\log{n}}{6n}}$, well above the edge deletion threshold. Unfortunately the blow-up operation also inflates independent sets. In [campos2025] it is shown that if one then runs a tailored nibble (loosely speaking, a guided triangle-free process), one will eventually double the density of the blow-up graph and regain the necessary pseudo-randomness, destroying the (structured) large independent sets. Our construction replaces the nibble in the second stage of [campos2025] by a second copy of a blow-up of the smaller random graph. We recover the necessary pseudorandom properties by randomly overlaying the two blow-ups; each destroys the large independent sets in the other.
@@
We note that a similar idea of stacking triangle-free graphs to reduce the independence number was used before by Alon and Rödl [Alon05] where they find bounds for multi-color Ramsey numbers $R(3,\ldots,3,k)$ by combining several copies of a Ramsey graph for $R(3,k)$. A second idea we utilize in our construction is that of sampling a subgraph from a larger graph to reduce the relative size of the independence number. Mubayi and Verstraëte use this idea in [MubayiVer24] to deduce lower bounds on Ramsey numbers from the existence of graphs with good pseudorandom properties (there measured by eigenvalues). A similar idea is used by Mattheus and Verstraëte [MattheusVer24] in their celebrated construction providing (up to logarithmic factors) a tight lower bound on $R(4,k)$.
@@
Let us now formally define our construction of the desired $n$ vertex graph $G$. For the sake of readability, we omit floors and ceilings whenever they are inconsequential. All of our logarithms are natural with base $e$. For any $n$ and $\varepsilon$, set
$N$ $=\frac{n}{\log^{2}(n)},$ $p$ $=\beta\sqrt{\frac{\log n}{n}},$ $k$ $=\kappa\sqrt{n\log n},$
where ultimately
$\beta=\frac{1}{2},$ $\kappa=1+\varepsilon.$
The construction then consists of the following steps:
@@
    Sample two (independent) copies of the random graph $G(N,p)$, $G_{R}$ and $G_{B}$, on vertex sets $V_{R}=\{r_{1},\dots,r_{N}\}$ and $V_{B}=\{b_{1},\dots,b_{N}\}$, with $p=\beta\sqrt{\frac{\log n}{n}}.$ We will refer to edges in $G_{R}$ and $G_{B}$ as red and blue edges respectively.
@@
    Let $G_{1}=G_{R}*G_{B}$ be the co-normal product, with vertex set $V_{R}\times V_{B}$ and $((r_{i},b_{j}),(r_{k},b_{\ell}))\in E(G_{1})$ if $(r_{i},r_{k})\in E(G_{R})$ or $(b_{j},b_{\ell})\in E(G_{B})$. See Figure 1 for a small example of such a product $G_{1}$. We call an edge $((r_{i},b_{j}),(r_{k},b_{\ell}))\in E(G_{1})$ red if $(r_{i},r_{k})\in E(G_{R})$ and blue if $(b_{j},b_{\ell})\in E(G_{B})$, noting that it is possible for a single edge to be colored both red and blue.
    $G_{B}$$G_{R}$..............................
    Figure 1: A small example of $G_{1}$
@@
    For any vertex $r_{i}\in V_{R}$, let $N(r_{i})$ denote the neighborhood of $r_{i}$ in $V_{R}$, and $N^{+}(r_{i}):=N(r_{i})\cap\{r_{i+1}\ldots,r_{N}\}$ the ‘right’ neighbors which lie above $r_{i}$ in the natural ordering of $V(G_{R})$. Denote the sets  $X_{r_{i}}=N(r_{i})\times V_{B}$ and $X_{r_{i}}^{+}=N^{+}(r_{i})\times V_{B}$; similarly define $X_{b_{i}}=V_{R}\times N(b_{i})$ and $X_{b_{i}}^{+}=V_{R}\times N^{+}(b_{i})$ for $b_{i}\in V_{B}$.
    We create a triangle-free graph $G_{2}\subseteq G_{1}$ by deleting all red edges in
    $\bigcup_{i=1}^{N}\left({X_{r_{i}}^{+}\choose 2}\cup{X_{b_{i}}\choose 2}\right)$,
    and all blue edges in
    $\bigcup_{i=1}^{N}\left({X_{r_{i}}\choose 2}\cup{X_{b_{i}}^{+}\choose 2}\right)$.
    Indeed, suppose that $(r_{i_{1}},b_{j_{1}}),(r_{i_{2}},b_{j_{2}}),(r_{i_{3}},b_{j_{3}})$ form a triangle in $G_{1}$. If this triangle is monochromatic red, then necessarily $r_{i_{1}},r_{i_{2}},r_{i_{3}}$ are distinct, and by relabeling we assume that $r_{i_{1}}<r_{i_{2}}<r_{i_{3}}$. Then the red edge $((r_{i_{2}},b_{j_{2}}),(r_{i_{3}},b_{j_{3}}))$ is removed as it belongs to $\binom{X_{r_{i_{1}}}^{+}}{2}$ (and an analogous argument holds for monochromatic blue triangles). Suppose then that $(r_{i_{1}},b_{j_{1}}),(r_{i_{2}},b_{j_{2}}),(r_{i_{3}},b_{j_{3}})$ is a triangle with two red edges and (after relabeling) $((r_{i_{1}},b_{j_{1}}),(r_{i_{2}},b_{j_{2}}))$ is blue. Then $((r_{i_{1}},b_{j_{1}}),(r_{i_{2}},b_{j_{2}}))\in\binom{X_{r_{3}}}{2}$ and this blue edge would be removed. The case of two blue edges and one red edge is identical.
@@
    Let $V(G)=\{v_{1},\ldots,v_{n}\}$, choose an injection $\varphi:V(G)\rightarrowtail V_{R}\times V_{B}$ uniformly at random, and include an edge $(v,v^{\prime})$ in $E(G)$ if and only if $(\varphi(v),\varphi(v^{\prime}))\in E(G_{2})$. It will be useful in the future to denote the graph induced by $G_{2}$ on $\varphi(V(G))$ as $G_{3}$ so that $G$ is the pre-image of $G_{3}$.
@@
    $0<\varepsilon_{2}\ll\varepsilon_{1}\ll\varepsilon\ll 1.$
    First we will use $\varepsilon_{2}$ in Lemma 3.1 to quantify standard concentration results for degrees and codegrees in random graphs, and $\varepsilon_{1}$ will be used in Lemmas 3.23.33.4, and 3.5 to quantify properties of $k$-sets in $G$. Finally in Section 4 we will show that $\alpha(G)<(1+\varepsilon)\sqrt{n\log n}$ with high probability.
3 Preliminaries
In this section, we introduce some notation and prove several structural lemmas that enable us to analyze the independent sets in $G$. For $r_{i}\in V_{R}$ let $F(r_{i})=\{(r_{i},b_{j})\in V(G_{3})\}$ and for $b_{i}\in V_{B}$ let $F(b_{i})=\{(r_{j},b_{i})\in V(G_{3})\}$ denote the respective fibers in $V(G_{3})$. Let
$N_{3}(r_{i})$ $=\bigcup_{r_{j}\in N(r_{i})}F(r_{j}),$ $N_{3}^{+}(r_{i})$ $=\bigcup_{r_{j}\in N^{+}(r_{i})}F(r_{j}),$
$N_{3}(b_{i})$ $=\bigcup_{b_{j}\in N(b_{i})}F(b_{j}),$ $N_{3}^{+}(b_{i})$ $=\bigcup_{b_{j}\in N^{+}(b_{i})}F(b_{j}).$
For any set $X\subseteq V_{R}\times V_{B}$ denote the projections onto $V_{R}$ and $V_{B}$ by
$\pi_{R}(X)$ $=\{r_{i}\in V_{R}:F(r_{i})\cap X\neq\emptyset\},$
$\pi_{B}(X)$ $=\{b_{i}\in V_{B}:F(b_{i})\cap X\neq\emptyset\}.$
For any set $X\subseteq V(G)$ we slightly abuse the notation, letting $\pi_{R}(X)=\pi_{R}(\varphi(X))$ and $\pi_{B}(X)=\pi_{B}(\varphi(X))$.

### v3

@@ Abstract
We present a flexible random construction which, for certain graphs $H$, is able to produce $H$-free graphs with edge density strictly larger than that of the $H$-free process, while simultaneously preserving pseudorandom properties and allowing a much easier analysis.
As our main application, we use this construction to show that the off-diagonal Ramsey numbers satisfy $R(3,k)\geq\left(\frac{1}{2}+o(1)\right)\frac{k^{2}}{\log{k}}$, improving the previously best bound $R(3,k)\geq\left(\frac{1}{3}+o(1)\right)\frac{k^{2}}{\log{k}}$. While the best known upper bound is $R(3,k)\leq\left(1+o(1)\right)\frac{k^{2}}{\log{k}}$, the constant of $\frac{1}{2}$ has been conjectured to be asymptotically tight by multiple groups.
11footnotetext: 2020 Mathematics Subject Classification: 05D10, 05D40, 05C55.
@@ 1 Introduction
The Ramsey number $R(\ell,k)$ is the least $N$ such that every $N$-vertex graph contains either a clique of size $\ell$ or an independent set of size $k$. In 1930 Ramsey [18] showed the existence of $R(\ell,k)$, and in the intervening century much effort has been dedicated to understanding the asymptotic behavior of these numbers. The best current bounds on the diagonal Ramsey numbers $R(k,k)$ (which are known exactly only up to $k=4$, with $43\leq R(5,5)\leq 46$ and the upper bound only very recently announced by Angeltveit and McKay [4]) are
@@
for some $c>0$. The lower bound is one of the earliest instances of Erdős’ probabilistic method [12], and the upper bound was the first exponential improvement over a result of Erdős and Szekeres [11], only recently announced in 2023 by Campos, Griffiths, Morris and Sahasrabudhe [6], and subsequently optimized to a value of $c\approx 0.2$ (for sufficiently large $k$) by Gupta, Ndiaye, Norin and Wei [15].
Perhaps the second most studied family of Ramsey numbers are the extreme off-diagonal numbers $R(\ell,k)$ for $\ell$ fixed and $k$ tending to infinity, where the case $\ell=3$ yields the first non-trivial asymptotic question. The exciting history of the study of $R(3,k)$ until the early 2000s is described by Spencer [21] in a chapter of the book “Ramsey Theory” [20], while the more recent history is very well captured by Campos, Jenssen, Michelen and Sahasrabudhe [8]. We briefly summarize the results determining the order of magnitude. The upper bound of $R(3,k)=O(\frac{k^{2}}{\log{k}})$ goes back to work of Ajtai, Komlós and Szemeŕedi  [1, 2] in the early 1980s, and in 1983 Shearer [19] reduced the implied constant to $R(3,k)\leq(1+o(1))\frac{k^{2}}{\log{k}}$. The lower bound was improved in several steps until Kim [16] proved in 1995 that $R(3,k)\geq c\frac{k^{2}}{\log{k}}$, with a constant $c\approx\frac{1}{160}$. Kim’s construction uses a probabilistic method commonly known as a “nibble”, where small batches of edges are added randomly for many steps. Doing so carefully, one can avoid the appearance of triangles while keeping the edge distribution sufficiently random to destroy large independent sets. Spencer [21] wrote in 2011:
@@
Nevertheless, only a short time later in 2013, two groups announced that they had greatly improved on the constant in the lower bound, showing that $R(3,k)\geq\left(\frac{1}{4}+o(1)\right)\frac{k^{2}}{\log{k}}$. Bohman and Keevash [5] and Fiz Pontiveros, Griffiths and Morris [13] very carefully analyzed the triangle-free process, in which an edge, chosen uniformly at random from those which would not create a triangle, is added to the graph (and this operation is repeated until no such edges exist). Both of these papers go through serious technical efforts to control the randomness in the process, and in particular show that with high probability the triangle-free process terminates with independence number yielding this bound. Fiz Pontiveros, Griffiths and Morris went so far as to conjecture that $c=\frac{1}{4}$ is the correct constant for $R(3,k)$.
In another breakthrough 12 years later, Campos, Jenssen, Michelen and Sahasrabudhe [8] showed that the story of the lower bound was not finished, and in fact $R(3,k)\geq\left(\frac{1}{3}+o(1)\right)\frac{k^{2}}{\log{k}}$. They conjecture that $c=\frac{1}{2}$ should be the correct constant.
Conjecture 1.1 (Campos, Jenssen, Michelen and Sahasrabudhe [8]).
@@
A conjecture by Davies, Jenssen, Perkins, and Roberts [10] relating maximum and average size of independence sets in triangle-free graphs would imply the upper bound in Conjecture 1.1. In this paper, we prove the lower bound.
@@
Similarly to all other previous proofs of lower bounds, we prove the following theorem concerning independence numbers, which easily implies Theorem 1.2.
@@
Spencer states in the epilogue of his book chapter:
@@
While inspired by ideas in [8], our proof of Theorem 1.3 does not require a nibble, and thus we are able to circumvent many of the technical difficulties faced by previous groups. If one were interested only in the correct order of magnitude, our construction could achieve Spencer’s dream, by combining slightly weaker ingredients with Shearer’s one page proof of the upper bound.
In Section 2 we present our random construction, which is quite versatile and has already seen applications to related problems. In Section 5 we give a direct application to hypergraph Ramsey numbers of stars, suggested to us by Mubayi. Campos, Jenssen, Michelen and Sahasrabudhe together with the fourth author [7] have built on our method to find the first polynomial improvements over the edge deletion threshold for the cycle-complete Ramsey numbers $r(C_{\ell},K_{k})$, for all odd $\ell\geq 9$ as $k$ tends to infinity. This shows some flexibility in the choice of an excluded graph $H$ beyond triangles. Most recently, Kühn, Sauermann, Steiner and Wigderson [kühn2025disproofoddhadwigerconjecture] have disproved the odd Hadwiger conjecture using our construction. Here, triangles are avoided the same way as in our application, guaranteeing a large chromatic number of the complement. Instead of avoiding large cliques in the complement, the authors show that one can avoid large odd pairings (see their paper for a definition), another pseudorandom property.
@@ 2 The Construction of $G$
Before formally presenting our construction witnessing Theorem 1.3, we provide some heuristic justification. The edge deletion method is a typical starting point in modern combinatorics when one wishes to construct a large graph $G$ which avoids a forbidden subgraph $H$, but carries certain other properties held by the random graph (in our application, $G$ should have no large independent sets). Implementing the method first requires one to calculate the edge deletion threshold $p_{H}$, so that (in expectation) the number of copies of $H$ in the random graph $G(n,p_{H})$ is of the same order as $p_{H}n^{2}$, the expected number of edges. If $p\ll p_{H}$ and one samples $G$ from $G(n,p)$ before deleting an edge from every copy of $H$, only $o(pn^{2})$ edges are lost.
Some further effort can show that the resulting graph will have small independence number (similar to $G(n,p)$) and many other pseudorandom properties. For a number of problems, finding constructions which improve upon this deletion method is of great interest. To outperform such examples, one may try to construct an $H$-free graph with edge density larger than $p_{H}$, while still being careful to obtain some of the pseudorandomness that the random graph enjoys.
@@
For triangles, the edge deletion threshold is of order $p_{K_{3}}=\sqrt{\frac{1}{n}}$. Most recently, Campos, Jenssen, Michelen and Sahasrabudhe [8] constructed a triangle-free graph with edge density $(1+o(1))\sqrt{\frac{2\log n}{3n}}$. To do so, they let ${m}=n/\log^{2}(n)$ and $p=\sqrt{\frac{\log(n)}{6n}}$ before sampling a graph $G_{1}$ from G(m,p)G({m},p). Noting that $p=o\left(\sqrt{\frac{1}{{m}}}\right)$, they can delete few edges to make the resulting graph (still on m{m} vertices) triangle-free, before blowing up each vertex by a factor $\log^{2}(n)$ to obtain a new graph on $n$ vertices, still triangle-free and with edge density $\sqrt{\frac{\log{n}}{6n}}$, well above the edge deletion threshold. Unfortunately the blow-up operation also inflates independent sets. In [8] it is shown that if one then runs a tailored nibble (loosely speaking, a nibble with an added regularization step), one will eventually double the density of the blow-up graph and destroy the (structured and rare) large independent sets created by the blow-up.
The main idea in our construction is to replace the nibble in the second stage of [8] by a second copy of a blow-up of the smaller random graph. We recover the necessary pseudorandom properties by randomly overlaying the two blow-ups; each copy of the blow-up destroys the large independent sets in the other. While there may be many resulting triangles from the overlay, they come in bunches intersecting in a single edge, and they can all be efficiently removed via the deletion of few edges.
@@
We learned that a similar idea of stacking triangle-free graphs to reduce the independence number was used before by Alon and Rödl [3] where they find lower bounds for multi-color Ramsey numbers $R(3,\ldots,3,k)$ by combining several copies of a Ramsey graph for $R(3,k)$. While the effect on the independence number is similar, they do not have to worry about triangles created from multiple colors.
In any graph $G$ witnessing a lower bound to $R(3,k)$, the neighborhood of a vertex $v$ is necessarily an independent set. The construction in [8] has independent sets asymptotically $\frac{3}{2}$ times the size of the average vertex degree. In our construction, independence number and average vertex degree asymptotically agree, and they are the same as a random graph of the same density. Any construction improving on the constant $\frac{1}{2}$ would have to have lower density, and at the same time independence number smaller than the random graph of that density.
@@
Let us now formally define our construction of the desired $n$ vertex graph $G$. For the sake of readability, we omit floors and ceilings whenever they are inconsequential. All of our logarithms are natural with base $e$. For some (to be specified) values of $p\in(0,1)$ and $s>1$, let ${m}=\frac{n}{s}$.
@@
    Sample two (independent) copies of the random graph G(m,p)G({m},p), $G_{R}$ and $G_{B}$ (we will call these the red and the blue graph) on vertex sets $V_{R}=V(G_{R})=\{r_{1},r_{2},\ldots,r_{{m}}\}$ and $V_{B}=V(G_{B})=\{b_{1},b_{2},\ldots,b_{{m}}\}$, respectively.
@@
    Let $V(G)=\{v_{1},v_{2}\ldots,v_{n}\}$, and choose a function $\pi:V(G)\to V_{R}\times V_{B}$ uniformly at random from all such injective maps. See below for a discussion of this choice.
@@
    Define functions $\pi_{R}:V(G)\to V_{R}$, $\pi_{B}:V(G)\to V_{B}$ by $(\pi_{R}(v_{i}),\pi_{B}(v_{i}))=\pi(v_{i})$ for all $i$. For any set $X\subseteq V(G)$, define the sets $\pi(X),\pi_{R}(X),\pi_{B}(X)$ as the images of the sets under these maps. We say that $\pi_{R}(X)$ and $\pi_{B}(X)$ are the projections of $\pi(X)$.
@@
    Define an edge set $\tilde{E}(G)=\tilde{E}_{R}(G)\cup\tilde{E}_{B}(G)$, where
    $\tilde{E}_{R}(G)$ $=\{v_{i}v_{j}:\pi_{R}(v_{i})\pi_{R}(v_{j})\in E(G_{R})\},$
    $\tilde{E}_{B}(G)$ $=\{v_{i}v_{j}:\pi_{B}(v_{i})\pi_{B}(v_{j})\in E(G_{B})\}.$
    If an edge is in $\tilde{E}_{R}(G)\cap\tilde{E}_{B}(G)$, we keep both edges, so $\tilde{E}(G)$ is a multigraph. Edges in $\tilde{E}_{R}(G)$ will be colored red, edges in $\tilde{E}_{B}(G)$ will be colored blue.
  5. 5.
    Order all pairs $r_{i}r_{j}\in{V_{R}\choose 2}$ lexicographically, i.e. first by the smaller index and then by the larger index. Similarly, order all pairs $b_{i}b_{j}\in{V_{B}\choose 2}$.
  6. 6.
    To arrive at $E(G)\subseteq\tilde{E}(G)$, we delete at least one edge from every triangle in $\tilde{E}(G)$.
    1. (a)
      From every red triangle $v_{i}v_{j}v_{\ell}\subseteq\tilde{E}_{R}(G)$, with $\pi_{R}(v_{j})\pi_{R}(v_{\ell})$ later in the order of the pairs in $V_{R}$ than $\pi_{R}(v_{i})\pi_{R}(v_{j})$ and $\pi_{R}(v_{i})\pi_{R}(v_{\ell})$, remove $v_{j}v_{\ell}\in\tilde{E}_{R}(G)$.
    2. (b)
      From every blue triangle $v_{i}v_{j}v_{\ell}\subseteq\tilde{E}_{B}(G)$ with $\pi_{B}(v_{j})\pi_{B}(v_{\ell})$ later in the order of the pairs in $V_{B}$ than $\pi_{B}(v_{i})\pi_{B}(v_{j})$ and $\pi_{B}(v_{i})\pi_{B}(v_{\ell})$, remove $v_{j}v_{\ell}\in\tilde{E}_{B}(G)$.
    3. (c)
      From every triangle $v_{i}v_{j}v_{\ell}$ with red edges $v_{i}v_{j},v_{i}v_{\ell}\in\tilde{E}_{R}(G)$, and blue edge $v_{j}v_{\ell}\in\tilde{E}_{B}(G)$, remove $v_{j}v_{\ell}\in\tilde{E}_{B}(G)$.
    4. (d)
      From every triangle $v_{i}v_{j}v_{\ell}$ with blue edges $v_{i}v_{j},v_{i}v_{\ell}\in\tilde{E}_{B}(G)$, and red edge $v_{j}v_{\ell}\in\tilde{E}_{R}(G)$, remove $v_{j}v_{\ell}\in\tilde{E}_{R}(G)$.
  7. 7.
    For our analysis, it is easier to work with this multigraph instead of the simple graph we obtain when we reduce every double edge to a single edge. As the independence number of the graph and the multigraph is the same, we will work with the multigraph.
The resulting multigraph on $E(G)=E_{R}(G)\cup E_{B}(G)$ is triangle-free and has density $(2+o(1))p$ for an appropriate choice of $\pi$, $s$ and $p$. Note that if every vertex in $V_{R}\cup V_{B}$ appears as $\pi_{R}(v_{i})$ or $\pi_{B}(v_{i})$ $(1+o(1))s$ times, then each of the edges removed in step 6 lies between the end vertices of at least $(1+o(1))s$ monochromatic paths of length $2$, and therefore every edge removal destroys at least $(1+o(1))s$ triangles in this process. This factor $s$ represents exactly the gain in efficiency of this construction when compared with the typical edge deletion method. A larger value of $s$ yields a larger gain, but the resulting graph is also further from a random graph.
There are many different ways to choose the function $\pi$ in step 2. Ultimately, it will be important that $\pi$ has certain expansion properties. In this paper, we like to think of $G$ as a random subgraph of the co-normal product of $G_{R}$ and $G_{B}$ on $V_{R}\times V_{B}$ to make the notion of a projection more natural, and we choose $\pi:V(G)\to V_{R}\times V_{B}$ as an injective function uniformly at random. In [kühn2025disproofoddhadwigerconjecture], the authors choose each $\pi(v_{i})$ independently uniformly at random, allowing repetition. In [7], the authors create the random union of balanced $s$-blow-ups of $G_{R}$ and $G_{B}$ alluded to in the heuristic description of our construction by choosing two random balanced partitions of the vertex set, which corresponds to choosing $\pi(v_{i})=(\pi_{R}(v_{i}),\pi_{B}(v_{i}))$ uniformly at random allowing repetition conditioned on each $r_{i}$ and $b_{i}$ being chosen exactly $s$ times by $\pi_{R}$ and $\pi_{B}$, respectively. Each of these choices have only marginal effects on the construction and the following analysis, and are largely interchangeable in each of the three applications.
For some applications it may be useful to replace the random choice in step 2 by an algebraic or geometric construction, e.g. the set of incidences of a projective plane. Overall, the construction is very flexible in the choices of $\pi$, $p$ and $s$ as long as one finds a variant of step 6 to efficiently destroy all copies of a given $H$.
We now define the parameters in the construction for our application to $R(3,k)$. For any $n$ and $\varepsilon$, set
$s$ $=\log^{2}(n),$ ${m}$ $=\frac{n}{s},$ $p$ $=\beta\sqrt{\frac{\log n}{n}},$ $k$ $=\kappa\sqrt{n\log n},$
where ultimately $\beta=\frac{1}{2}$ and $\kappa=1+\varepsilon$.
@@
$0<\varepsilon_{2}\ll\varepsilon_{1}\ll\varepsilon\ll 1,$
and in addition, $n_{0}=n_{0}(\varepsilon_{2})\gg\frac{1}{\varepsilon_{2}}$ is taken to be sufficiently large so that certain inequalities hold for all $n>n_{0}$. We will always treat $\varepsilon,\varepsilon_{1},\varepsilon_{2}$ as fixed, and $o(1)$ notation refers to the asymptotics as $n$ grows.
In Section 3, we will use $\varepsilon_{2}$ in Lemma 3.1 to quantify routine concentration results in our construction, and $\varepsilon_{1}$ in Lemmas 3.23.33.4, and 3.5 to control the quantity and location of the removed edges in $\tilde{E}(G)\setminus E(G)$.
Finally in Section 4 we will show that $\alpha(G)<(1+\varepsilon)\sqrt{n\log n}$ with high probability. For this, we will look at the projections of a given $k$-set onto $G_{R}$ and $G_{B}$. If we disregard the deleted edges, a $k$-set in $G$ is independent if and only if both the projections in $G_{R}$ and $G_{B}$ are independent. Thus, a $k$-set with large projections is less likely to be independent than a set with small projections. On the other hand, there are many more $k$-sets with large projections than with small projections. A careful balance of these counts and probabilities together with the information gathered on the removed edges in Section 3 then enables us to prove the desired bound.
Throughout, we use the standard notation $N(v)$ for the neighborhood of a vertex in a graph, where the graph to consider is often implied by the choice of the vertex $v$.
3 Controlling the edge deletions
In this section, we introduce some notation and prove several structural lemmas that enable us to analyze the $k$-sets in $G$ and help us to control the deleted edges in $\tilde{E}(G)\setminus E(G)$. For $r_{i}\in V_{R}$ and $b_{i}\in V_{B}$ let
$F(r_{i})$ $=\pi_{R}^{-1}(r_{i}),$ $F(b_{i})$ $=\pi_{B}^{-1}(b_{i})$
denote the respective fibers in $V(G)$.
The edges in $\tilde{E}_{R}$ incident to a vertex $v_{i}\in V(G)$ are independent of $\pi_{B}$ (and symmetrically for $\tilde{E}_{B}$ and $\pi_{R}$). In fact, if $\pi_{R}(v_{i})=\pi_{R}(v_{j})$, then the red neighborhoods under $\tilde{E}_{R}$ of $v_{i}$ and $v_{j}$ are identical. We can describe red and blue neighborhoods in $\tilde{E}$ as follows (where the $N_{v}^{+}$ will play a role when we consider edge deletions in monochromatic triangles):
$N_{r_{i}}$ $=\bigcup_{r_{j}\in N(r_{i})}F(r_{j}),$ $N_{r_{i}}^{+}$ $=\bigcup_{j>i,r_{j}\in N(r_{i})}F(r_{j}),$
$N_{b_{i}}$ $=\bigcup_{b_{j}\in N(b_{i})}F(b_{j}),$ $N_{b_{i}}^{+}$ $=\bigcup_{j>i,b_{j}\in N(b_{i})}F(b_{j}).$